### v1

@@ 2.1 Complexity analysis
Theorem 2.1 shows that the iteration complexity of Algorithm 1 for finding a $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$. This is consistent with the best iteration complexity of the second-order algorithm Luo2022FindingSS ; Chen2023ACR ; Yao2024TwoTR for solving non-convex-strongly concave minimax problems, omitting the logarithmic factor, and is better than the best iteration complexity of the first-order algorithm. In Yang2023AccI , a first-order algorithm is proposed to obtain the second-order stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\kappa^{1.75}\epsilon^{-1.75})$. Compared with this algorithm, Algorithm 1 converges slightly faster, but each step requires solving the gradient regularization trust region subproblem.
Moreover, we can have a rough estimate of the number of inner iterations, i.e., $N_{t}$. If $x_{t}$ is not an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point, according to Lemmas 2.5 and 2.6, we conclude that either $\|s_{t}\|=r\max\{\|g_{t}\|^{1/2},\epsilon^{1/2}\}$ or $\|g_{t}\|\leqslant(L_{1}+\sigma\|g_{t}\|^{1/2})\|s_{t}\|$ for the case $\lambda_{t}=0$ and $\|g_{t}\|\geqslant\epsilon$. Then it follows that
$\|s_{t}\|\geqslant\mathcal{O}(\min\{\sqrt{\frac{\epsilon}{L_{2}}},\frac{\epsilon}{L_{1}+\sqrt{L_{2}}}\}).$
By further combining with Lemma 2.4, we obtain that $A=\mathcal{O}(\rho^{0.5}\kappa^{0.5}\ell^{-2}\epsilon^{1.5}+\rho^{-0.5}\kappa^{-0.5}\epsilon)$, $N_{1}=\mathcal{O}(\kappa\log\frac{\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\kappa\log\frac{A+\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2$). If the Nesterov’s accelerated gradient algorithm is used to solve the subproblem regarding $y$ as in Luo2022FindingSS , then the number of inner iterations will reduce to $N_{1}=\mathcal{O}(\sqrt{\kappa}\log\frac{\sqrt{\kappa+1}\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\sqrt{\kappa}\log(\sqrt{\kappa+1}+\frac{\sqrt{\kappa+1}\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2$) by Theorem 2 in Luo2022FindingSS .
@@ 3 A Levenberg-Marquardt algorithm with negative curvature correction
In the classic nonlinear minimization optimization, when the Hessian matrix is not positive definite, the negative curvature direction can be used to correct the iteration direction of the algorithm to converge to a second-order stationary point Gratton2023YetAF . In this section, we further propose a Levenberg-Marquardt (LM) algorithm with a gradient norm regularization coefficient Mischenko2023RegularizedNM , which uses the negative curvature direction to correct the iteration direction (LMNegCur) to solve (1.1) without solving the trust region subproblem at each iteration. The LMNegCur algorithm consists of the following two important algorithmic components at each iteration:
  • Gradient ascent update: for given $x_{t}$, run $N_{t}$ iterations of gradient ascent step to obtain an approximated maximizer
@@
  • LM or negative curvature update: for a given pair $(x_{t},y_{t})$, if $\lambda_{\min}(H_{t})\leqslant-\frac{\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}}{2}$, perform a negative curvature update
    $s_{t}=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}u_{t},~{}\text{where}~{}g_{t}^{\top}u_{t}\leqslant 0,\|u_{t}\|=1,H_{t}u_{t}=\lambda_{\min}(H_{t})u_{t};$
@@
Note that for the case $\|g_{t}\|\leqslant\epsilon,~{}\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}$, we have already achieved an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$, which will be proved later. The smallest eigenvalue of $H_{t}$ and the corresponding curvature direction can be retrieved via a Lanczos procedure. Since $\|H_{t}\|\leqslant L_{1}$, it was shown in Royer2018ComplexityAO (Lemma 9) that after at most $n$ (the dimension of $x$) iterations, the procedure obtains a unit vector $u$ such that $u^{\top}H_{t}u=\lambda_{\min}(H_{t})$ with probability 1. Therefore, using the smallest eigenvalue of the Hessian does not incur much cost.
@@
  Step 1: Input $x_{0},y_{0},\eta_{y},N_{t},\epsilon$; Set $t=1$, $x_{1}=x_{0}$.
@@
   (a): Set $k=0$, $y_{k}^{t}=y_{t-1}$.
   (b): Update $y_{k}^{t}$:
$y_{k+1}^{t}=y_{k}^{t}+\eta_{y}\nabla_{y}f(x_{t},y_{k}^{t}).$ (3.1)
@@
  Step 3: If $\lambda_{\min}(H_{t})\leqslant-\frac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$,
$s_{t}=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}u_{t},~{}\text{where}~{}g_{t}^{\top}u_{t}\leqslant 0,\|u_{t}\|=1,H_{t}u_{t}=\lambda_{\min}(H_{t})u_{t};$ (3.2)
      otherwise, go to Step 4.
  Step 4: If $\|g_{t}\|\geqslant\epsilon$,
$s_{t}=-(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)^{-1}g_{t};$ (3.3)
@@
  Step 5: Update $x_{t}$: $x_{t+1}=x_{t}+s_{t}.$ Set $t=t+1,$ go to Step 2(a).

### v2

@@ 2.1 Complexity analysis
Theorem 2.1 shows that the outer iteration of Algorithm 1 for finding an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$. This is consistent with the best known iteration complexity of the second-order algorithm Chen2023ACR ; Luo2022FindingSS ; Yao2024TwoTR for solving nonconvex-strongly concave minimax problems, omitting the logarithmic factor, and is better than the best iteration complexity of the first-order algorithm. In Yang2023AccI , a first-order algorithm is proposed to obtain the second-order stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$. Compared with this algorithm, Algorithm 1 converges slightly faster, but each step requires solving the gradient regularization trust-region subproblem. Theorem 2.1 also implies that the total number of gradient ascent steps required is $\tilde{\mathcal{O}}(\ell^{2}\rho^{0.5}\mu^{-2}\epsilon^{-1.5})$, which matches the algorithm in Chen2023ACR ; Luo2022FindingSS .
2.2 Inexact gradient norm regularized trust-region algorithm
It should be noted that the complexity analysis in Section 2.1 is based on the exact solution of the trust-region subproblem (2.3). However, in practice, the trust-region subproblem cannot always be solved exactly. Moreover, when dealing with large-scale problems, the computational cost of obtaining such exact solutions can become very high, especially when matrix decomposition and Hessian-vector products are involved. In this section, we propose an Inexact Gradient Norm Regularized Trust-Region (IGRTR) algorithm that allows subproblems to be solved inexactly and can be proven to have the same iteration complexity as the GRTR algorithm.
Unlike the GRTR algorithm, when updating $x$, the IGRTR algorithm approximately solves the trust-region subproblem (2.3) based on the following two cases:
  • When $\|g_{t}\|>\epsilon$, it solves the trust-region subproblem (2.3) in the same way as the trust-region Newton-CG method (Algorithm 4.1 in Curtis2021TrustN ), obtaining a direction $s_{t}$ that satisfies either of the following conditions:
    $g_{t}^{\top}s_{t}+\dfrac{1}{2}s_{t}^{\top}H_{t}s_{t}\leqslant-\dfrac{1}{8}\sigma\|g_{t}\|^{1/2}\|s_{t}\|^{2},~{}\|s_{t}\|=r\|g_{t}\|^{1/2},$ (2.34)
    or
    $\begin{cases}g_{t}^{\top}s_{t}+\dfrac{1}{2}s_{t}^{\top}H_{t}s_{t}\leqslant-\dfrac{1}{8}\sigma\|g_{t}\|^{1/2}\|s_{t}\|^{2},~{}\|s_{t}\|<r\|g_{t}\|^{1/2},\\ \|(H_{t}+\sigma\|g_{t}\|^{1/2}I)s_{t}+g_{t}\|\leqslant\dfrac{1}{2}\min\bigg{\{}\|g_{t}\|,\dfrac{1}{2}\sigma\|g_{t}\|^{1/2}\|s_{t}\|\bigg{\}};\end{cases}$ (2.35)
  • When $\|g_{t}\|\leqslant\epsilon$, the algorithm applies the randomized Lanczos procedure (Algorithm 3.2 in Curtis2021TrustN , Kuczynski1992Estimating ) to $H_{t}$, either $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sigma\epsilon^{1/2}$, or else obtains a direction $s_{t}$ that satisfies
    $g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}\leqslant-\frac{1}{8}\sigma\epsilon^{1/2}\|s_{t}\|^{2},~{}\|s_{t}\|=r\epsilon^{1/2}.$ (2.36)
Note that from Lemmas 2.4 and (2.1), it is easy to verify that an exact solution to the trust-region subproblem (2.3) satisfies these conditions. By replacing Step 4 in the GRTR algorithm with the above procedure, the detailed IGRTR algorithm for solving (1.1) can be immediately obtained.
In the next section, we will show that in order to solve (1.1) for an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$, the outer iteration complexity of the IGRTR algorithm is also $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$, and the required Hessian-vector product is at most $\tilde{\mathcal{O}}(\ell^{0.5}\rho^{0.25}\kappa^{1.75}\epsilon^{-1.75})$.
2.3 Complexity analysis
In the following lemmas, we give key inequalities on the descent of the function value and the norm of the gradient for two different cases $\|g_{t}\|\geqslant\epsilon$ and $\|g_{t}\|\leqslant\epsilon$.
Lemma 2.9
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be the sequence generated by the IGRTR algorithm. Set
$\sigma=\frac{8\sqrt{L_{2}}}{9},r=\frac{1}{3\sqrt{L_{2}}},\epsilon_{1}\leqslant\min\bigg{\{}\frac{1}{162},\frac{\sqrt{L_{2}}}{162L_{1}}\bigg{\}}\epsilon^{3/2},\epsilon_{2}\leqslant\frac{\sqrt{L_{2}}}{27}\epsilon^{1/2},$ (2.37)
and choose $\eta_{y},\theta,N_{t}$ the same as in Lemma 2.4. For the case $\|g_{t}\|>\epsilon$, if $\|s_{t}\|=r\|g_{t}\|^{1/2}$, we have
$P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{486\sqrt{L_{2}}}\|g_{t}\|^{3/2},~{}\|g_{t+1}\|\leqslant 2+\frac{L_{1}}{3\sqrt{L_{2}}}\|g_{t}\|^{1/2}+\|g_{t}\|;$ (2.38)
if $\|s_{t}\|<r\|g_{t}\|^{1/2}$, we have
$P(x_{t+1})\leqslant P(x_{t}),~{}\|g_{t+1}\|\leqslant\frac{1}{2}\|g_{t}\|.$ (2.39)
Proof
According to (2.1), (2.34) and (2.35), we can get
$P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{8}\sigma\|g_{t}\|^{1/2}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (2.40)
Substituting (2.37) and $\|s_{t}\|=r\|g_{t}\|^{1/2}$ into the above inequality, we get the first inequality of (2.38). The second inequality of (2.38) is the same as that in (2.1).
Next, we consider the case where $\|s_{t}\|<r\|g_{t}\|^{1/2}$. To bound the right-hand side of (2.40), we first give some bounds on $\|g_{t}\|$. By combining (2.35) with $\|H_{t}\|\leqslant\|H_{t}-\nabla^{2}P(x_{t})\|+\|\nabla^{2}P(x_{t})\|\leqslant L_{1}+\epsilon_{2}$, we can easily obtain
$\|g_{t}\|\leqslant$ $~{}\|(H_{t}+\sigma\|g_{t}\|^{1/2}I)s_{t}+g_{t}\|+\|(H_{t}+\sigma\|g_{t}\|^{1/2}I)s_{t}\|$
$\leqslant$ $~{}\frac{1}{2}\|g_{t}\|+\big{(}L_{1}+\epsilon_{2}+\sigma\|g_{t}\|^{1/2}\big{)}\|s_{t}\|.$
This means
$\epsilon^{3/2}\leqslant\epsilon^{1/2}\|g_{t}\|\leqslant 2(L_{1}+\sigma)\|g_{t}\|^{1/2}\|s_{t}\|+2\epsilon_{2}\|s_{t}\|.$ (2.41)
The rest of the proof is the same as Lemma 2.5. We omit it here for simplicity.
Lemma 2.10
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be the sequence generated by the IGRTR algorithm, with all parameters chosen the same as in Lemma 2.9. Then, for the case of $\|g_{t}\|\leqslant\epsilon$, if $\lambda_{\min}(H_{t})\geqslant-\frac{\sigma}{2}\epsilon^{1/2}$, then we have found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$; otherwise, we have
$P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{486\sqrt{L_{2}}}\epsilon^{3/2},~{}\|g_{t+1}\|\leqslant 2+\frac{L_{1}}{3\sqrt{L_{2}}}\epsilon^{1/2}.$ (2.42)
Combined with (2.36), the following proof is basically the same as the proof of Lemma 2.6 and will not be repeated here.
Let $\epsilon>0$ be any given target accuracy. We denote the first iteration of the IGRTR algorithm to obtain $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\frac{4}{9}\sqrt{L_{2}\epsilon}$ as
$\hat{T}(\epsilon):=\min\left\{t|\|g_{t}\|\leqslant\epsilon~{}\text{and}~{}\lambda_{\min}(H_{t})\geqslant-\dfrac{4}{9}\sqrt{L_{2}\epsilon}\right\}.$
Then, by Lemma 2.10, we know that $x_{\hat{T}(\epsilon)}$ is already an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point.
$\mathcal{F}_{\hat{T}(\epsilon)}=\left\{t<\hat{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{486\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\right\},$ (2.43)
$\mathcal{G}_{\hat{T}(\epsilon)}=\left\{t<\hat{T}(\epsilon):\|g_{t+1}\|\leqslant\frac{1}{2}\|g_{t}\|,t\notin\mathcal{F}_{\hat{T}(\epsilon)}\right\}.$ (2.44)
From Lemmas 2.9 and 2.10, we conclude that each iteration $t$ belongs to at least one of the above two sets $\mathcal{F}_{\hat{T}(\epsilon)}$ and $\mathcal{G}_{\hat{T}(\epsilon)}$.
Similar to the proof of Theorem 2.1 in Section 2, we give an upper bound on $\hat{T}(\epsilon)$ in the following theorem.
Theorem 2.2
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by the IGRTR algorithm with all parameters chosen the same as in Lemma 2.9. We have
$\hat{T}(\epsilon)\leqslant$ $~{}\hat{C}_{1}\epsilon^{-3/2}\ln\epsilon^{-1}+\hat{C}_{2}\epsilon^{-3/2}+1=\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5}),$
$\sum_{t=0}^{\hat{T}(\epsilon)}N_{t}\leqslant 2\sqrt{\kappa}\ln$ $\frac{\sqrt{\kappa+1}\|y_{-1}-y^{*}(x_{0})\|}{A}+2\sqrt{\kappa}\hat{T}(\epsilon)\ln\frac{\sqrt{\kappa+1}(A+\kappa r\sqrt{\hat{G}})}{A},$
where $\hat{C}_{1}=486\sqrt{L_{2}}(P(x_{0})-P^{*})/\ln 2$, $\hat{C}_{2}=(\ln\hat{G}+\ln 2)\hat{C}_{1}$, $\hat{G}:=\max\big{\{}2+\frac{L_{1}}{3\sqrt{L_{2}}}\hat{C}^{1/2}+\hat{C},2+\frac{L_{1}}{3\sqrt{L_{2}}}\big{\}}$ with $\hat{C}:=\left(486\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}$, and $A$ is defined as in Lemma 2.4.
Next, we analyze the Hessian-vector product required by the IGRTR algorithm.
Corollary 2.1
Suppose that Assumption 2.1 holds. The matrix inverse computation $(\nabla_{yy}^{2}f(x_{t},y_{t}))^{-1}$ in $H_{t}$ can be approximated by $\tilde{\mathcal{O}}(\sqrt{\kappa})$ terms of Chebyshev polynomials as described in Luo2022FindingSS . Let $\{x_{t}\}$ be the sequence generated by the IGRTR algorithm, and all other parameters are chosen the same as in Lemma 2.9. Then it can be shown that $x_{\hat{T}(\epsilon)}$ is an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with high probability guarantee. Therefore, the total number of Hessian-vector products required is at most $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$.
Proof
Before analyzing the Hessian vector products required by the IGRTR algorithm, we first show that replacing $(\nabla_{yy}^{2}f(x_{t},y_{t}))^{-1}$ with an approximation does not affect the iteration complexity bound in Theorem 2.2. Denote
$\hat{H}_{t}:=\nabla_{xx}^{2}f(x_{t},y_{t})+\nabla_{xy}^{2}f(x_{t},y_{t})C_{t}\nabla_{yx}^{2}f(x_{t},y_{t})~{}\text{with}~{}C_{t}\approx-(\nabla_{yy}^{2}f(x_{t},y_{t}))^{-1}.$
By the choice of $N_{t}$ in Lemma 2.4 and Lemma 8 in Luo2022FindingSS , it can be proved that
$\|\hat{H}_{t}-\nabla^{2}P(x_{t})\|\leqslant\|\hat{H}_{t}-H_{t}\|+\|H_{t}-\nabla^{2}P(x_{t})\|\leqslant\epsilon_{2},$
This shows that the iteration complexity result in Theorem 2.2 is not affected when $\hat{H}_{t}$ is used instead of $H_{t}$. In addition, for any vector $u$, we can obtain $\hat{H}_{t}u$ by calling $\tilde{\mathcal{O}}(\sqrt{\kappa})$ Hessian-vector products (Appendix E Luo2022FindingSS ).
Note that by Theorem 2.2, the outer iteration complexity is $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$. In each outer iteration, according to Lemmas 3.1 and 4.1 in Curtis2021TrustN and Lemma 2 in Royer2020ANewton , approximately solving the trust-region subproblem (2.3) requires at most $\min\big{\{}n,\tilde{\mathcal{O}}(L_{1}^{0.5}L_{2}^{-0.25}\epsilon^{-0.25})\big{\}}$ iterations to obtain a direction $s_{t}$ that satisfies conditions (2.34), (2.35), and (2.36) with probability $1-\xi$, where $\xi$ is the failure probability of the Lanczos procedure. Since the failure probability $\xi$ is only within a ”logarithmic factor” of the complexity bound, we can use a very small $\xi=\hat{\xi}/(\hat{C}_{1}\epsilon^{-3/2}\ln\epsilon^{-1}+\hat{C}_{2}\epsilon^{-3/2}+1)$, where $\hat{\xi}\in(0,1)$, $\hat{C}_{1}$ and $\hat{C}_{2}$ are defined in Theorem 2.2. Therefore, $x_{\hat{T}(\epsilon)}$ is an expected second-order stationary point with probability $1-\hat{\xi}$, and the total number of Hessian-vector products required is
$\tilde{\mathcal{O}}(\sqrt{\kappa})\hat{T}(\epsilon)\min\big{\{}n,\tilde{\mathcal{O}}(L_{1}^{0.5}L_{2}^{-0.25}\epsilon^{-0.25})\big{\}}\leqslant\tilde{\mathcal{O}}(\ell^{0.5}\rho^{0.25}\kappa^{1.75}\epsilon^{-1.75}).$
The proof is complete.
Remark 2.2
Compared with Theorem 2.1, the above Theorem 2.2 and Corollary 2.1 show that the number of outer iterations and gradient ascent iterations of the IGRTR algorithm are consistent with the order of the GRTR algorithm. Compared to the IMCN algorithm Luo2022FindingSS and the ICLM algorithm Chen2023ACR , both of which require $\tilde{\mathcal{O}}(\ell^{2.5}\mu^{-1.5}\epsilon^{-2})$ Hessian-vector products, the IGRTR algorithm only requires $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$ Hessian-vector products, which has a better complexity bound in terms of the order of $\epsilon$.
@@ 3 A Levenberg-Marquardt algorithm with negative curvature correction
In the classic nonlinear minimization optimization, when the Hessian matrix is not positive definite, the negative curvature direction can be used to correct the iteration direction of the algorithm to converge to a second-order stationary point Gratton2023YetAF . In this section, we further propose a Levenberg-Marquardt (LM) algorithm with a gradient norm regularization coefficient Mischenko2023RegularizedNM , which uses the negative curvature direction to correct the iteration direction (LMNegCur) to solve (1.1) without solving the trust-region subproblem at each iteration. The LMNegCur algorithm consists of the following two important algorithmic components at each iteration:
  • Gradient ascent update: for given $x_{t}$, for a given $x_{t}$, run $N_{t}$ iterations of Nesterov’s accelerated gradient ascent step to obtain an approximated maximizer
@@
  • LM or negative curvature update: for a given pair $(x_{t},y_{t})$, if $\lambda_{\min}(H_{t})\leqslant-\frac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, perform a negative curvature update
    $s_{t}=\sqrt{\max\{\|g_{t}\|,\epsilon\}/L_{2}}u_{t},~{}\text{where}~{}g_{t}^{\top}u_{t}\leqslant 0,\|u_{t}\|=1,~{}\lambda_{\min}(H_{t})=u_{t}^{\top}H_{t}u_{t};$
@@
  Step 1: Input $x_{0},y_{-1},\eta_{y},\theta,N_{t},\epsilon$; Set $t=0$.
@@
   (a): Set $k=0$, $y_{k}^{t}=\tilde{y}_{k}^{t}=y_{t-1}$.
   (b): Update $y_{k}^{t}$ and $\tilde{y}_{k}^{t}$:
$y_{k+1}^{t}=\tilde{y}_{k}^{t}+\eta_{y}\nabla_{y}f(x_{t},\tilde{y}_{k}^{t}),~{}\tilde{y}_{k+1}^{t}=y_{k+1}^{t}+\theta(y_{k+1}^{t}-y_{k}^{t}).$ (3.1)
@@
  Step 3: Set $g_{t}=\nabla_{x}f(x_{t},y_{t})$ and $H_{t}=H(x_{t},y_{t})$.
  Step 4: Compute an eigenpair $(\lambda_{\min}(H_{t}),u_{t})$ such that
$\lambda_{\min}(H_{t})=u_{t}^{\top}H_{t}u_{t},\|u_{t}\|=1,~{}g_{t}^{\top}u_{t}\leqslant 0.$ (3.2)
     If $\lambda_{\min}(H_{t})\leqslant-\frac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$,
$s_{t}=\sqrt{\max\{\|g_{t}\|,\epsilon\}/L_{2}}u_{t};$ (3.3)
     otherwise, go to Step 5.
  Step 5: If $\|g_{t}\|\geqslant\epsilon$,
$(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)s_{t}=-g_{t};$ (3.4)
@@
  Step 6: Update $x_{t}$: $x_{t+1}=x_{t}+s_{t}.$ Set $t=t+1,$ go to Step 2(a).
Note that for the case $\|g_{t}\|\leqslant\epsilon,~{}\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}$, we have already achieved an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$, which will be proved later. Differing from the trust-region type Yao2024TwoTR and cubic regularization type algorithms Luo2022FindingSS ; Chen2023ACR , the LMNegCur algorithm is free of further inner iterative processes in the involved outer iteration and only requires at most one negative curvature computation and one solution of a positive definite linear system. In addition, the hard-case of trust-region subproblem and cubic regularization subproblem, in which the gradient $g_{t}$ is orthogonal to the eigenspace of the Hessian $H_{t}$ corresponding to its minimum eigenvalue, does not present a serious challenge. The minimum eigenvalue of $H_{t}$ and the corresponding curvature direction can be retrieved via a Lanczos procedure Kuczynski1992Estimating . Since $\|H_{t}\|\leqslant L_{1}$, it was shown in Royer2018ComplexityAO (Lemma $9$) that after at most $n$ (the dimension of $x$) iterations, the procedure obtains a unit vector $u$ such that $u^{\top}H_{t}u=\lambda_{\min}(H_{t})$ with probability $1$. As also noted in Chapter $10$ of Golub2013 Matrix , for large and sparse problem, information about the extremal eigenvalues of matrix tends to emerge quite early in the iteration of the Lanczos procedure. Then, using the smallest eigenvalue of the Hessian does not incur much cost in such case.